You have two capacitors, one with capacitance 10.5×10^−6 F and the other of unknown capacitance. You connect the two capacitors in series with a voltage of 335 V applied across the capacitor pair. You discover that, as a result, the unknown capacitor has a charge of 0.00199 C . Find its capacitance ? .
Answer #1

Solution,

V1 = Q1/C1 = 0.00199/10.5 x 10^-6 = 189.52 V

V2 = 335 - 189.52 = 145.48 V

C2 = Q2/V2 = 0.00199/145.48 = 13.68 uF

Comment in case any doubt please rate my answer.. .